# Searching for Structure: Investigating Emergent Communication with Large Language Models

### v1

@@ 3.3 LLMs as agents
The LLMs in our experiment were instruction-tuned instantiations of Llama 3 70B Llama Team (2024) with greedy sampling222Although we only report results on one model type, initial explorations with GPT-3.5 and Llama 2 7B showed similar behaviours to LLama 3 70B. While human participants typically learn signal-meaning mappings through a learning block, we use LLMs’ in-context learning Brown et al. (2020) ability to teach them the languages. Specifically, we prepend our prompts with the items to be learned in a structured JSON-like format (Prompt 1). Given the observed behavioural similarities between humans and LLMs Galke et al. (2023), we assume that a vocabulary of signal-meaning mappings in the context of a prompt provides enough (distributional) information for a LLM to learn an appropriate mapping between the attributes of the stimuli and signal syllables. Although the prompt structure ’invites’ the LLM to infer a signal from the stimulus attributes, we are agnostic about how exactly and what kind of mapping the LLM deduces, but we are interested in the resulting behaviours.
@@
Generating signals. For signal generation in the labelling block, we use prompt completion (Prompt 1). During labelling, the agents see the entire training set and generate a signal for each stimulus, effectively amounting to a look-up task since the stimulus is present in the prompt. On the other hand, the vocabulary presented to agents during communication and testing does not include the current stimulus, thus requiring the agents to extract an appropriate mapping and generalise to new stimuli (Prompt 2). A human-like solution would be to map stimulus attributes (i.e. shape, colour, and amount) to syllables representing these attributes and create compositions that describe the stimulus. During communication, we add a communicativeSuccess attribute which is set to if the previous interaction for this stimulus was successful and zero otherwise. Adding this attribute functions as a memory between interactions and provides a pressure for expressivity. It is hypothesised that the latter plays an important role in human language evolution since it prevents languages from becoming degenerate Smith et al. (2013). Importantly, during testing, the vocabulary presented to the agents always includes the train set (without the current stimulus), and items from the test set are never present.
Guessing signals or meanings. For guessing and discrimination during communication, the agents need to respond with a choice corresponding to the speaker’s signal. Unfortunately, LLMs are inconsistent and unreliable in answering multiple choice questions Khatun and Brown (2024). In our initial exploration, this indeed proved to be unusable. Instead, for each distractor (signal or meaning), we run the prompt prefilled with that distractor through the model and select the distractor with the highest probability (Prompt 3). Again, the agents observe the training vocabulary with the current stimulus in the guessing block. In the communication block, agents observe the training vocabulary without the current stimulus.
@@ 3.4 Metrics
We are firstly interested in investigating whether two agents settle on a language that enables them to communicate, measured by the percentage of successful interactions (PercCom) in a round. We use multiple metrics to measure structure in messages. The most common metric is topographic similarity (TopSim, Brighton and Kirby, 2006). Similar to Kirby et al. (2008), we report Z-scores of the Mantel test Mantel (1967) between signal similarities (normalised Levenshteins distance) and semantic similarities (the number of equal attributes between two meanings). A communication system with a high TopSim uses similar signals for similar meanings. We compute the Ngram diversity Meister et al. (2023), being the average fraction of unique vs. total Ngrams for in all produced signals. Low Ngram diversity implies the agents re-use parts of signals in different signals, hinting at compositional signals when it happens in combination with increased TopSim. We assess the degree of signal systematicity between the signals produced for unseen stimuli in the test block and the previous stimuli in the communication block using the generalisation score (GenScore, Raviv et al., 2021). First, we compute the pairwise semantic difference between each stimulus in the train and test scenes, followed by the pairwise normalised edit distance between the signals produced for these scenes. We then take the Pearson correlation between these differences across all stimuli. Intuitively, this measures whether similar scenes across both sets are similarly labelled, thereby suggesting generalisation.
@@ 5.2 Agents communicate successfully
Once the agents have individually learned the vocabulary, they start communicating. Despite initially starting with different languages, of the interactions in the first round are successful (chance performance would amount to ). This increases somewhat in the following rounds to , but not significantly (Appendix A, Figure 5). Interestingly, communicative success is not guaranteed, it fluctuates between rounds and in some simulations it even decreases drastically.
@@ 7 Discussion
The rapid increase in learnability resulting from iterated learning proves that weak learning biases in language models, such as, for example, an observed simplicity bias Chen et al. (2024), can be amplified by the process of generational transmission. Additional simulations with increased communicative difficulty, e.g., by increasing the number of distractors or the number of interaction partners, could reveal whether and how some form of memory constraint affects the learnability of the languages. Doing so additionally captures the diversity and dynamic nature of language in the real world. In general, systematic manipulations across model features (e.g., size, training data, or decoding strategies) may expose why we observe tendencies such as producing longer signals. Similar to what was proposed by Galke and Raviv (2024), we argue that careful manipulation of our setup can help reveal underlying mechanistic biases of language models and inform modelling choices when simulating language acquisition in LLMs. Taking into account the important role communication plays in shaping human language, LLM performance drastically increased when it was optimised for successful communication through reinforcement learning from human feedback.
@@ Appendix C Prompts
Our agents act based on prompts and system instructions. These are designed to be maximally close to the classical experimental setup formatted similar to Galke et al. (2023). The completion prompt 1 is used for labelling and guessing. For the latter task, we prefill the word and pick the word with the highest probability. See Prompts 1, 2, and 3 below.

### v2

@@ 3.3 LLMs as agents
The LLMs in our experiment were instruction-tuned instantiations of Llama 3 70B Llama Team (2024) with greedy sampling222Although we only report results on one model type, initial explorations with GPT-3.5 and Llama 2 7B showed similar behaviours to LLama 3 70B. While human participants typically learn signal-meaning mappings through a learning block, we use LLMs’ in-context learning Brown et al. (2020) ability to teach them the languages. Specifically, we prepend our prompts with the items to be learned in a structured JSON-like format (Prompt 1). Given the observed behavioural similarities between humans and LLMs Galke et al. (2023), we assume that a vocabulary of signal-meaning mappings in the context of a prompt provides enough (distributional) information for a LLM to learn an appropriate mapping between the attributes of the stimuli and signal syllables. Although the prompt structure ‘invites’ the LLM to infer a signal from the stimulus attributes, we are agnostic about how exactly and what kind of mapping the LLM deduces, but we are interested in the resulting behaviours.
@@
Generating signals. For signal generation in the labelling block, we use prompt completion (Prompt 2). During labelling, the agents see the entire training set and generate a signal for each stimulus, effectively amounting to a look-up task since the stimulus is present in the prompt. On the other hand, the vocabulary presented to agents during communication and testing does not include the current stimulus, thus requiring the agents to extract an appropriate mapping and generalise to new stimuli (Prompt 3). A human-like solution would be to map stimulus attributes (i.e. shape, colour, and amount) to syllables representing these attributes and create compositions that describe the stimulus. During communication, we add a communicativeSuccess attribute which is set to if the previous interaction for this stimulus was successful and zero otherwise. Adding this attribute functions as a memory between interactions and provides a pressure for expressivity. It is hypothesised that the latter plays an important role in human language evolution since it prevents languages from becoming degenerate Smith et al. (2013). Importantly, during testing, the vocabulary presented to the agents always includes the train set (without the current stimulus), and items from the test set are never present.
Guessing signals or meanings. For guessing and discrimination during communication, the agents need to respond with a choice corresponding to the speaker’s signal. Unfortunately, LLMs are inconsistent and unreliable in answering multiple choice questions Khatun and Brown (2024). In our initial exploration, this indeed proved to be unusable. Instead, for each distractor (signal or meaning), we run the prompt prefilled with that distractor through the model and select the distractor with the highest probability (Prompt 4). Again, the agents observe the training vocabulary with the current stimulus in the guessing block. In the communication block, agents observe the training vocabulary without the current stimulus.
@@ 3.4 Metrics
We are firstly interested in investigating whether two agents settle on a language that enables them to communicate, measured by the percentage of successful interactions (PercCom) in a round. We use multiple metrics to measure structure in messages. The most common metric is topographic similarity (TopSim, Brighton and Kirby, 2006). Similar to Kirby et al. (2008), we report Z-scores of the Mantel test Mantel (1967) between signal similarities (normalised Levenshteins distance) and semantic similarities (the number of equal attributes between two meanings). A communication system with a high TopSim uses similar signals for similar meanings. We compute the Ngram diversity Meister et al. (2023), being the average fraction of unique vs. total Ngrams for in all produced signals. Low Ngram diversity across all signals implies the agents re-use parts of signals in different signals, hinting at compositional signals when it happens in combination with increased TopSim. We assess the degree of signal systematicity between the signals produced for unseen stimuli in the test block and the previous stimuli in the communication block using the generalisation score (GenScore, Raviv et al., 2021). Here, we first compute the pairwise semantic difference between each stimulus in the train and test scenes, followed by the pairwise normalised edit distance between the signals produced for these scenes. We then take the Pearson correlation between these differences across all stimuli. Intuitively, this measures whether similar scenes across both sets are similarly labelled, thereby suggesting generalisation.
@@ 5.2 Agents communicate successfully
Once the agents have individually learned the vocabulary, they start communicating. Despite initially starting with different languages, approximately of the interactions in the first round are successful (chance performance would amount to ). This increases somewhat in the following rounds to , but not significantly (Appendix A, Figure 5). Interestingly, communicative success is not guaranteed, it fluctuates between rounds and in some simulations it even decreases drastically.
@@ 7 Discussion
The rapid increase in learnability resulting from iterated learning proves that weak learning biases in language models, such as, for example, an observed simplicity bias Chen et al. (2024), can be amplified by the process of generational transmission. Additional simulations with increased communicative difficulty, e.g., by increasing the number of distractors or the number of interaction partners, could reveal whether and how some form of memory constraint affects the learnability of the languages. Doing so additionally captures the diversity and dynamic nature of language in the real world. In general, systematic manipulations across model features (e.g., size, training data, or decoding strategies) may expose why we observe tendencies such as producing longer signals. Similar to what was proposed by Galke and Raviv (2024), we argue that careful manipulation of our setup can help reveal underlying mechanistic biases of language models and inform modelling choices when simulating language acquisition in LLMs. Taking into account the important role communication plays in shaping human language, LLM performance drastically increased when it was optimised for successful communication through reinforcement learning from human feedback (RLHF).
@@ Appendix C Prompts
Our agents act based on prompts and system instructions. These are designed to be maximally close to the classical experimental setup and formatted similar to Galke et al. (2023). The completion prompt 2 is used for labelling and guessing. For the guessing task, we prefill the word and pick the signal with the highest probability. See the full prompts for labelling and guessing (Prompt 2), speaking (Prompt 3), discrimination (Prompt 4) below.